y=-x-9 find the y intercept
Answer by moshiz08(60)

The equation is in the form y=mx+b, where b tells you the y-intercept. In this case b=-9.
The y-intercept is defined to be the point at which the graph intersects the y-axis. That occurs when the x-coordinate is zero. Plug in x=0 to the equation y=-x-9 to get y=-0-9 = -9. So the y-intercept will be the point (0, -9).